# FedEFC: Federated Learning Using Enhanced Forward Correction Against Noisy Labels

### v1

@@ 3.3 Objective Function Analysis in Noisy Labels
We analyze the objective function of FL in noisy labels when applying FedEFC. Each participating client updates its model by minimizing the local loss function during each round. Clients optimize their models iteratively. When using stochastic gradient descent (SGD), the update rule is given by ${w}_{k}^{t+1}\leftarrow{w}_{k}^{t}-\eta_{k}\nabla\mathcal{L}_{k}(\bm{x},y;{w}_{k}^{t})$ where $\eta_{k}$ denotes the learning rate of client $k$. According to [13], the objective function of FL is formulated as
@@
The objective function in FL with noisy labels can be formulated using a composite loss, which combines a proper loss with a link function. A proper loss is used for class probability estimation, while a link function maps the classifier’s output to the range $[0,1]$. A composite loss is a proper loss, as proven in [17]. In the FL system, local loss function corresponds to a proper loss and the inverse of the softmax function acts as the link function ($\sigma^{-1}:[0,1]\rightarrow\mathbb{R}$). Since composite loss is defined as the combination of a proper loss and a link function, the composite loss for client $k$ is expressed as
$(\mathcal{L}_{k})^{\sigma^{-1}_{k}}({y},{\bm{h}^{k}(\bm{x}};w))=\mathcal{L}_{k}({y},\sigma_{k}({\bm{h}^{k}(\bm{x};w}))).$ (18)
In the case of a clean dataset, the minimizer of the loss function, following the property of composite proper losses, is given by
$\arg\min_{\bm{h}}\mathbb{E}_{\bm{x},{y}}[(\mathcal{L}_{k})^{\sigma^{-1}_{k}}({y},\bm{h}(\bm{x};w))]$ (19)
$=\sigma_{k}^{-1}(p({y}|\bm{x};w)).$ (20)
Therefore, the FL objective function can be rewritten as
$\arg\min_{w}{F}_{g}({w})$
$=\sum_{k\in\mathcal{S}_{t}}\frac{n_{k}\arg\min_{\bm{h}}\mathbb{E}_{\bm{x},{y}}[(\mathcal{L}_{k})^{\sigma^{-1}_{k}}({y},\bm{h}(\bm{x};w))]}{\sum_{u\in\mathcal{S}_{t}}n_{u}}.$ (21)
For a noisy dataset, the minimizer of the loss function is replaced by $\sigma_{k}^{-1}(p(\tilde{y}|\bm{x};w))$. However, under FedEFC, the objective function evaluated at the minimizer for clean data remains identical to that evaluated at the minimizer for noisy data when using enhanced forward correction. Theorem 1 demonstrates this equivalence, as derived in [16].
@@
Let every $Q_{{\tilde{y}}|{{y}}}^{k}$ generated by client $k$ be a non-singular matrix. A composite loss incorporating $Q^{k}_{{\tilde{y}}|{y}}$ is given by
$(\mathcal{L}_{k}^{\textbf{F}})^{\sigma^{-1}_{k}}({y},{\bm{h}^{k}(\bm{x}};w))$ $=\mathcal{L}_{k}({y},Q^{k}_{{\tilde{y}}|{y}}\cdot\sigma_{k}({\bm{h}^{k}(\bm{x};w})))$
Then, the objective function evaluated at the minimizer for clean data is equivalent to that evaluated at the minimizer for noisy data:
$\sum_{k\in\mathcal{S}_{t}}\frac{n_{k}\arg\min_{\bm{h}}\mathbb{E}_{\bm{x},{\tilde{y}}}[(\mathcal{L}_{k}^{\textbf{F}})^{\sigma^{-1}_{k}}({y},\bm{h}(\bm{x};w))]}{\sum_{u\in\mathcal{S}_{t}}n_{u}}$
$=\sum_{k\in\mathcal{S}_{t}}\frac{n_{k}\arg\min_{\bm{h}}\mathbb{E}_{\bm{x},{y}}[(\mathcal{L}_{k})^{\sigma^{-1}_{k}}({y},\bm{h}(\bm{x};w))]}{\sum_{u\in\mathcal{S}_{t}}n_{u}}$
Proof. : We analyze the objective function by examining the minimizer of the loss function for client $k$. The minimizer of the loss function using the enhanced forward correction is derived as follows using $\phi_{k}=Q^{k}_{{\tilde{y}}|{y}}\cdot\sigma_{k}$:
$\arg\min_{\bm{h}}\mathbb{E}_{\bm{x},{\tilde{y}}}[(\mathcal{L}_{k}^{\textbf{F}})^{\sigma^{-1}_{k}}({y},\bm{h}(\bm{x};w))]$
$=\arg\min_{\bm{h}}\mathbb{E}_{\bm{x},{\tilde{y}}}[(\mathcal{L}_{k})^{\phi^{-1}_{k}}({y},\bm{h}(\bm{x};w))]$ (22)
$=\phi^{-1}_{k}(p({\tilde{y}}|\bm{x}))$ (23)
$=\sigma_{k}^{-1}\left((Q^{k}_{{\tilde{y}}|{y}})^{-1}\cdot p({\tilde{y}}|\bm{x})\right)$ (24)
$=\sigma_{k}^{-1}\left((Q^{k}_{{\tilde{y}}|{y}})^{-1}\cdot Q^{k}_{{\tilde{y}}|{y}}\cdot p({{y}}|\bm{x})\right)$ (25)
$=\sigma_{k}^{-1}\left(I\cdot p({{y}}|\bm{x})\right)$ (26)
$=\arg\min_{\bm{h}}\mathbb{E}_{\bm{x},{y}}[(\mathcal{L}_{k})^{\sigma^{-1}_{k}}({y},\bm{h}(\bm{x};w))].$ (27)
For client $k$, the minimizer of the loss function using enhanced forward correction under noisy labels is shown to be equivalent to the minimizer of the loss function under clean labels. Since the summation of the minimizers forms the objective function, Theorem 1 holds.
@@ 4 Experiments
In this section, FedEFC is evaluated on non-IID data with noisy labels using the MNIST [8], CIFAR-10 [7], and CIFAR-100 [14] datasets. To assess robustness against noisy labels, noise parameters $\rho$ and $\zeta$ are varied under fixed heterogeneity. Additionally, experiments are conducted with various heterogeneous parameters $\alpha_{\text{dir}}$ and $p$ while keeping noise levels constant to simulate environments ranging from nearly IID to highly non-IID. By combining these parameters, diverse scenarios with varying $(\alpha_{\text{dir}},p)$ and $(\rho,\zeta)$ are constructed. We compare FedEFC with standard FL techniques, noise-robust FL approach, and CL for DNNs trained with noisy labels. The results demonstrate that FedEFC is robust against both noisy labels and data heterogeneity.
@@ 4.1 Experiment Setup
Non-IID with Noisy Labels : In a real-world scenario, noise is first introduced into the training dataset before it is allocated to all clients in the FL system. Since similar noisy labels occur across clients, a consistent noise pattern is applied before partitioning the dataset according to non-IID parameters $(\alpha_{\text{dir}},p)$. This procedure creates a non-IID environment in which noisy labels are class-dependent.
Baselines : The performance of FedEFC is compared with several FL techniques: FedAvg [13], FedDitto [11], FedDyn [1], and FedProx [10]. FedCorr [22], a noise-robust FL method designed to tackle noisy labels, is included. Furthermore, once the proposed prestopping point is determined, both forward correction [16] and confident learning [15] are applied in the FL setting for comparison. As a reference, FedAvg without noise is employed to validate the robustness of the noise by comparing performance against the ideal noise-free case.
@@ 4.2 Implementation Details
FL experiments are simulated with $N=100$ clients. Each client performs local training using SGD with a momentum of $0.5$ and at each round, a fraction $0.1$ of clients are randomly selected for aggregation. The dataset-specific parameters are summarized in Table 1. The prestopping threshold, $\gamma_{\text{thr}}$, is set to $3$ for MNIST and $6$ for CIFAR-10 and CIFAR-100. To ensure training stability, prestopping monitoring begins after round $10$ for MNIST and after round $40$ for CIFAR-10 and CIFAR-100.
In cases of severe label imbalance, particularly when $p=0.2$, estimating the noise transition matrix directly from the raw count matrix is challenging. To compensate for this issue, FedEFC refines the count matrix by weighting it with the class distribution of each client before normalizing its columns. This adjustment is applied to MNIST and CIFAR-10 when $p=0.2$, while the standard count matrix remains effective under other settings.
Table 1: Dataset and FL parameter settings.
Parameter MNIST CIFAR-10 CIFAR-100
Class Size 10 10 100
Number of Clients 100 100 100
Model 9-Layers CNN ResNet-18 ResNet-34
Rounds 100 200 300
Local Batch Size 64 10 10
Table 2: Average test accuracy and standard deviation over 3 trials under noise-fixed settings with noise amount $\rho=0.2$ and sparsity $\zeta=0.8$ for various non-IID parameter configurations. The best two values in each column are highlighted in bold, except for the noiseless case (FedAvg wo noise).
@@
confident learning $\bm{98.97\pm 0.13}$ $\bm{98.94\pm 0.16}$ $91.44\pm 2.75$ $\bm{88.01\pm 0.24}$ $80.21\pm 3.29$ $62.45\pm 1.21$ ${57.92\pm 0.11}$ $56.68\pm 0.83$ $47.49\pm 1.08$
forward correction ${97.05\pm 0.52}$ $94.79\pm 0.09$ $\bm{97.00\pm 0.20}$ $82.76\pm 0.21$ $\bm{80.62\pm 0.70}$ $\bm{70.80\pm 1.59}$  $56.72\pm 0.58$ $56.73\pm 0.22$ $\bm{52.00\pm 0.94}$
FedEFC $\bm{98.15\pm 0.45}$ $\bm{98.51\pm 0.46}$ $\bm{96.78\pm 0.60}$ $\bm{88.37\pm 0.08}$ $\bm{86.78\pm 0.36}$ $\bm{70.57\pm 1.25}$ $\bm{60.57\pm 0.40}$ $\bm{59.34\pm 0.51}$ $\bm{50.58\pm 0.18}$
FedAvg wo noise $99.49\pm 0.01$ $99.51\pm 0.02$ $98.91\pm 0.21$ $90.74\pm 0.14$ $88.93\pm 0.46$ $80.43\pm 0.86$ $68.14\pm 0.51$ $67.42\pm 0.32$ $63.58\pm 0.20$
Table 3: Average test accuracy and standard deviation over 3 trials under non-IID fixed settings with Dirichlet $\alpha_{\text{dir}}=10.0$ and Bernoulli $p=0.5$ for various noise configurations.
@@
FedCorr $65.71\pm 0.49$ $97.36\pm 0.60$ $\bm{99.19\pm 0.13}$ $54.06\pm 27.01$ $78.10\pm 1.42$ $80.31\pm 2.65$ $\bm{44.86\pm 0.07}$ $\bm{58.13\pm 1.57}$ $64.50\pm 1.78$
confident learning $\bm{68.38\pm 1.75}$ $\bm{99.22\pm 0.04}$ $\bm{99.33\pm 0.01}$ $\bm{66.20\pm 1.40}$ $\bm{84.95\pm 0.22}$ $\bm{86.83\pm 0.26}$ $41.37\pm 1.26$ $56.19\pm 0.62$ $64.51\pm 0.24$

### v3

@@ 3.3 Objective Function Analysis in Noisy Labels
We analyze the objective function of FL in noisy labels when applying FedEFC. Each participating client updates its model by minimizing the local loss function during each round. Clients optimize their models iteratively. When using stochastic gradient descent (SGD), the update rule is given by ${w}_{k}^{t+1}\leftarrow{w}_{k}^{t}-\eta_{k}\nabla\mathcal{L}_{k}(\bm{x},y;{w}_{k}^{t})$ where $\eta_{k}$ denotes the learning rate of client $k$. According to [14], the objective of FL is formulated as the minimization problem: Compute the value of
@@
The objective function in FL with noisy labels can be formulated using a composite loss, which combines a proper loss with a link function. A proper loss is used for class probability estimation, while a link function maps the classifier’s output to the range $[0,1]$. A composite loss is a proper loss, as proven in [18]. In the FL system, local loss function corresponds to a proper loss and the inverse of the softmax function acts as the link function ($\sigma_{k}^{-1}:[0,1]\rightarrow\mathbb{R}$). Since composite loss is defined as the combination of a proper loss and a link function, the composite loss $(\mathcal{L}_{k})^{\sigma^{-1}_{k}}$ for client $k$ is defined by $(\mathcal{L}_{k})^{\sigma^{-1}_{k}}({y},{\bm{h}^{k}(\bm{x}};w))=\mathcal{L}_{k}(\bm{x},y;w)$. In the case of a clean dataset, the minimizer of the loss function, following the property of composite proper losses, is given by
$\!\!\!\!\!\!\arg\min_{\bm{h}}\mathbb{E}_{\bm{x},{y}}[(\mathcal{L}_{k})^{\sigma^{-1}_{k}}({y},\bm{h}(\bm{x};w))]=\sigma_{k}^{-1}(p({y}|\bm{x};w)).$ (14)
Therefore, the objective of FL can be reformulated as the following minimization problem: Compute the value of
$\arg\min_{w}\Bigg(\sum_{k\in\mathcal{S}_{t}}\frac{n_{k}\arg\min_{\bm{h}}\mathbb{E}_{\bm{x},{y}}[(\mathcal{L}_{k})^{\sigma^{-1}_{k}}({y},\bm{h}(\bm{x};w))]}{\sum_{u\in\mathcal{S}_{t}}n_{u}}\Bigg).$ (15)
For a noisy dataset, the minimizer of the composite loss $(\mathcal{L}_{k})^{\sigma^{-1}_{k}}$ is instead given by $\sigma_{k}^{-1}(p(\tilde{y}|\bm{x};w))$. However, under FedEFC, the objective evaluated at the minimizer for clean data remains approximately equal to that for noisy data when enhanced forward correction is applied. Thm. 1 shows this approximate equivalence, as derived in [17].
@@
Assume each matrix $Q^{k}_{\tilde{y}|y}$ generated by client $k$ is non-singular, and approximately equal to the true transition matrix $T^{k}_{\tilde{y}|y}$ whose $(i,j)$th entries are true conditional probabilities $p(\tilde{y}=i|y=j)$. A composite loss incorporating $Q^{k}_{{\tilde{y}}|{y}}$ is given by
$(\mathcal{L}_{k}^{\mathbf{F}})^{\sigma^{-1}_{k}}(y,\bm{h}^{k}(\bm{x};w))=\mathcal{L}_{k}^{\mathbf{F}}(\bm{x},y;w).$ (16)
Then, the aggregated objective at the minimizer for clean data is approximately equal to that at the minimizer for noisy data:
$\sum_{k\in\mathcal{S}_{t}}\frac{n_{k}\,\arg\min_{\bm{h}}\mathbb{E}_{\bm{x},\tilde{y}}\bigl[(\mathcal{L}_{k}^{\mathbf{F}})^{\sigma^{-1}_{k}}(y,\bm{h}(\bm{x};w))\bigr]}{\sum_{u\in\mathcal{S}_{t}}n_{u}}$ (17)
$\;\;\approx\;\sum_{k\in\mathcal{S}_{t}}\frac{n_{k}\,\arg\min_{\bm{h}}\mathbb{E}_{\bm{x},y}\bigl[(\mathcal{L}_{k})^{\sigma^{-1}_{k}}(y,\bm{h}(\bm{x};w))\bigr]}{\sum_{u\in\mathcal{S}_{t}}n_{u}}.$
Proof. Fix client $k$ and define $\phi_{k}=Q^{k}_{{\tilde{y}}|{y}}\cdot\sigma_{k}$. Then
$\arg\min_{\bm{h}}\mathbb{E}_{\bm{x},\tilde{y}}\bigl[(\mathcal{L}_{k}^{\mathbf{F}})^{\sigma^{-1}_{k}}(y,\bm{h}(\bm{x};w))\bigr]$ (18)
$=\arg\min_{\bm{h}}\mathbb{E}_{\bm{x},\tilde{y}}\bigl[(\mathcal{L}_{k})^{\phi^{-1}_{k}}(y,\bm{h}(\bm{x};w))\bigr]$ (19)
$=\phi_{k}^{-1}(p(\tilde{y}|\bm{x}))=\sigma_{k}^{-1}\!\left((Q^{k}_{{\tilde{y}}|{y}})^{-1}p(\tilde{y}|\bm{x})\right)$ (20)
$=\sigma_{k}^{-1}\!\left((Q^{k}_{{\tilde{y}}|{y}})^{-1}T^{k}_{{\tilde{y}}|{y}}\,p(y|\bm{x})\right)\approx\sigma_{k}^{-1}\!\left(I\cdot p(y|\bm{x})\right)$ (21)
$=\arg\min_{\bm{h}}\mathbb{E}_{\bm{x},y}\bigl[(\mathcal{L}_{k})^{\sigma^{-1}_{k}}(y,\bm{h}(\bm{x};w))\bigr].$ (22)
For client $k$, the minimizer of the loss function using enhanced forward correction under noisy labels is shown to be approximately equivalent to the minimizer of the loss function under clean labels. Since the summation of the minimizers forms the objective function, Thm. 1 holds.
Remark. The accuracy of the approximation depends on how close $Q^{k}_{{\tilde{y}}|{y}}$ is to $T^{k}_{{\tilde{y}}|{y}}$.
@@ 4 Experiments
In this section, FedEFC is evaluated on non-IID data with noisy labels using the MNIST [8], CIFAR-10 [7], and CIFAR-100 [15] datasets. To assess robustness against noisy labels, we vary noise parameters $(\rho,\zeta)$ with fixed heterogeneity and, conversely, vary heterogeneity parameters $(\alpha_{\text{dir}},p)$ under fixed noise levels. This design enables simulation of environments from nearly IID to highly non-IID. By combining these parameters, diverse scenarios with varying $(\alpha_{\text{dir}},p)$ and $(\rho,\zeta)$ are constructed. We compare FedEFC with standard FL techniques, noise-robust FL approach, and CL-based methods originally developed for DNNs trained with noisy labels but extended to the FL setting. The results demonstrate that FedEFC is robust against both noisy labels and data heterogeneity.
@@ 4.1 Experiment Setup
Non-IID with Noisy Labels: To emulate a real-world scenario, noise is first introduced into the training dataset before it is allocated to all clients in the FL system. Since similar noisy labels occur across clients, a consistent noise pattern is applied before partitioning the dataset according to non-IID parameters $(\alpha_{\text{dir}},p)$. This procedure creates a non-IID environment in which noisy labels are class-dependent.
Baselines: We compare FedEFC with standard FL methods (FedAvg [14], FedDitto [10], FedProx [11]), noise-robust baselines (FedCorr [24], FedNoRo [23], RoFL [25]), and, once the prestopping point is determined, loss correction (forward correction [17]) and data pruning (confident learning [16]) methods adapted to FL. As a reference, FedAvg without noise is reported as the noise-free upper bound.
@@ 4.2 Implementation Details
FL experiments are simulated with $N=100$ clients. Each client performs local training using SGD with a momentum of $0.5$ and at each round, a fraction $0.1$ of clients are randomly selected for aggregation. For MNIST, we use a 9-layer CNN trained for $100$ rounds with batch size $64$; for CIFAR-10, ResNet-18 for $200$ rounds with batch size $10$; and for CIFAR-100, ResNet-34 for $300$ rounds with batch size $10$. The prestopping threshold, $\gamma_{\text{thr}}$, is empirically set to $3$ for MNIST and $6$ for CIFAR-10 and CIFAR-100. To ensure training stability, prestopping monitoring begins after round $10$ for MNIST and after round $40$ for CIFAR-10 and CIFAR-100.
In cases of severe label imbalance, particularly when $p=0.2$, estimating the noise transition matrix directly from the raw count matrix is challenging. To compensate for this issue, we apply a weighted noise transition matrix defined as $\widetilde{Q}^{k}_{\tilde{y}=i\mid y=j}=\frac{p^{k}_{i}\,\bm{C}^{k}_{\tilde{y}=i,y=j}}{\sum_{j}p^{k}_{j}\,\bm{C}^{k}_{\tilde{y}=i,y=j}}$ where $p^{k}_{i}=\frac{|\mathbf{X}_{\tilde{y}^{k}=i}|}{\sum_{j}|\mathbf{X}_{\tilde{y}^{k}=j}|}$. This weighting scheme compensates for label imbalance by incorporating each client’s class distribution. The adjustment is applied to MNIST and CIFAR-10 when $p=0.2$; otherwise, the standard count matrix remains effective.
Table 1: Average test accuracy ($\pm$ std over 3 trials) under noise-fixed settings with $\rho=0.2$, $\zeta=0.8$, across different non-IID parameters. The best two values per column are in bold, except for the noiseless case (FedAvg wo noise).
@@
RoFL $\bm{99.00\pm 0.05}$ $\bm{98.98\pm 0.05}$ $\bm{97.36\pm 0.50}$ $\bm{89.22\pm 0.16}$ $\bm{86.11\pm 0.13}$ $54.65\pm 0.31$ $58.55\pm 0.04$ $51.33\pm 0.45$ $40.97\pm 0.85$
FedNoRo $97.43\pm 0.13$ $95.56\pm 1.15$ $65.08\pm 4.62$ $84.19\pm 0.25$ $76.68\pm 2.44$ $37.54\pm 2.30$ $58.57\pm 1.27$} $43.07\pm 0.64$ $14.57\pm 1.94$
confident learning $\bm{98.97\pm 0.13}$ $\bm{98.94\pm 0.16}$ $91.44\pm 2.75$ $88.01\pm 0.24$} $80.21\pm 3.29$ $62.45\pm 1.21$ $57.92\pm 0.11$} $56.68\pm 0.83$ $47.49\pm 1.08$
forward correction $97.05\pm 0.52$} $94.79\pm 0.09$ $\bm{97.00\pm 0.20}$ $82.76\pm 0.21$ $80.62\pm 0.70$} $\bm{70.80\pm 1.59}$ $56.72\pm 0.58$ $56.73\pm 0.22$ $\bm{52.00\pm 0.94}$
FedEFC $98.15\pm 0.45$} $98.51\pm 0.46$} $96.78\pm 0.60$} $\bm{88.37\pm 0.08}$ $\bm{86.78\pm 0.36}$ $\bm{70.57\pm 1.25}$ $\bm{60.57\pm 0.40}$ $\bm{59.34\pm 0.51}$ $\bm{50.58\pm 0.18}$
FedAvg wo noise $9.49\pm 0.01 99$$9.51\pm 0.02$ $98.91\pm 0.21$ $90.74\pm 0.14$ $88.93\pm 0.46$ $80.43\pm 0.86$ $68.14\pm 0.51$ $67.42\pm 0.32$ $63.58\pm 0.20$
Table 2: Average test accuracy ($\pm$ std over 3 trials) under non-IID fixed settings $(\alpha_{\text{dir}},p)=(10.0,0.5)$ with varying noise configurations.
@@
[…]
RoFL $68.50\pm 0.27$ $\bm{99.21\pm 0.05}$ $99.14\pm 0.01$} $62.29\pm 2.53$ $\bm{86.91\pm 0.42}$ $\bm{87.85\pm 0.14}$ $34.94\pm 0.16$ $2.83\pm 0.21 55$$5.46\pm 0.05$
FedNoRo $\bm{70.17\pm 1.88}$ $97.23\pm 0.26$ $98.73\pm 0.06$ $58.94\pm 2.15$ $78.59\pm 0.70$ $83.47\pm 0.15$ $37.63\pm 0.64$ $43.38\pm 2.50$ $50.63\pm 1.43$
confident learning $68.38\pm 1.75$} $\bm{99.22\pm 0.04}$ $\bm{99.33\pm 0.01}$ $\bm{66.20\pm 1.40}$ $84.95\pm 0.22$} $86.83\pm 0.26$} $41.37\pm 1.26$ $56.19\pm 0.62$ $64.51\pm 0.24$
@@
Table 3: CIFAR-10 test accuracy ($\pm$ std, 3 runs) under high noise with $\rho=0.8$, $\zeta\in\{0.0,0.4,0.8\}$, and $(\alpha_{\text{dir}},p)=(100.0,0.8)$.
($\rho,\zeta)$ (0.8, 0.0) (0.8, 0.4) (0.8, 0.8)
FedAvg $21.88\pm 0.48$ $17.30\pm 0.28$ $18.75\pm 0.66$
FedDitto $21.97\pm 0.36$ $14.28\pm 0.58$ $19.38\pm 0.86$
FedProx $21.33\pm 0.33$ $17.14\pm 0.10$ $18.58\pm 0.33$
FedCorr $20.54\pm 0.67$ $14.30\pm 0.39$ $9.54\pm 1.31$
RoFL $25.38\pm 1.18$ $15.51\pm 0.21$ $17.52\pm 0.32$
FedNoRo $\bm{25.92\pm 1.82}$ $\bm{20.31\pm 0.04}$ $\bm{27.70\pm 1.22}$
confident learning $9.25\pm 0.53 11$$1.48\pm 0.29$ $18.49\pm 0.12$
forward correction $23.40\pm 0.31$ $17.02\pm 0.25$ $18.61\pm 0.24$
FedEFC $\bm{27.37\pm 1.55}$ $\bm{20.69\pm 3.94}$ $\bm{29.90\pm 3.06}$